# remarkRemark

### v1

@@ Abstract
A technique for determining this augmentation parameter is proposed, and the loss-reduction and convergence properties of the bSQH algorithm are analysed theoretically and validated numerically. Results of numerical experiments in the context of image classification with a sparsity enforcing $L^{0}$-based regularizer demonstrate the effectiveness of the proposed method in full-batch and mini-batch modes.
@@ 1 Introduction
In the last decades, convolutional neural networks (CNNs) have become a standard choice for computer vision tasks and, in particular, medical imaging analysis, where they are applied to aid in detecting and diagnosing diseases and classifying tumour types based on MRI, CT or X-ray scans. Compared to its predecessors, like the multilayer perceptron (MLP), CNNs utilize convolutional layers with trainable filters in addition to fully connected layers, making them especially suited for processing data with a grid-like topology. This network architecture enables a more efficient extraction of local patterns and relevant features from an input image, by simultaneously requiring fewer parameters than fully connected architectures. The development of CNNs started in the 1980s when Kunihiko Fukushima [1] introduced the ’Neocognitron’ as the earliest precursor of convolutional networks. Almost one decade later Yann LeCun [2] developed the ’LeNet-5’ architecture for document recognition, representing a milestone in the design of modern CNNs. Over the last decades, a variety of architectures have contributed to the success of these network types, amongst which there is AlexNet [3], which popularized the use of ReLU activation functions and VGGNet [4] where an increased network depth was used, showing the benefits of using smaller filter sizes. However, it occurs that very deep architectures often face a vanishing gradient problem, preventing very deep networks from being trained efficiently using gradient-based training methods and backpropagation. To mitigate this vanishing gradient problem, Kaiming He et al. [5] developed the ResNet, a network architecture that utilizes skip connections, which allows the efficient training of even deeper networks.
@@
Apart from the possibility of training deep networks, the introduction of ResNet-like architectures gave birth to the dynamical system viewpoint of deep learning, where networks with a residual structure are related to numerical schemes for solving differential equations; see [6, 7, 8, 9, 10, 11]. The dynamical systems viewpoint offers important insights into the stability of a network’s forward propagation process by allowing the study of very deep networks from a continuous-time dynamical systems perspective. More importantly for our investigation, it motivates the development of novel learning techniques based on optimal control theory [12, 13].
@@
While serving as a motivation for developing new training algorithms, the continuous-time dynamical systems interpretation can be limiting, particularly when considering network architectures without skip connections such as LeNet-5 and AlexNet. In those cases, the continuous-time interpretation becomes even more difficult when taking into account that CNN models use dimensionality reduction as a design approach for feature extraction. However, even if the relation to the continuous-time setting is lost for most modern network architectures, it has been shown that special optimal control-based techniques can still be applied to the machine learning framework, almost independently of the network’s structure [14]. Based thereon, in this work, we considerably extend our optimal control-based learning technique presented in [13] to the framework of CNNs. For this purpose, we construct a CNN as a discrete-time controlled system including trainable filters, dimensionality reduction in the form of pooling operations, and a sequence of fully-connected layers for classification. Further for this extension and within this special structure, we also take into account the computational complexity arising from the increased number of parameters and dimensionality of data, which can impose memory constraints. To address these challenges, we derive a novel mini-batch variant of our optimal control-based learning algorithm. Furthermore, we include sparse network training in the discussion of our algorithm, where the aim is to determine the essential parameters for a given model by forcing irrelevant or redundant ones to be zero.
@@
In the framework of CNNs, where models are often over-parametrized, deriving sparse solutions can greatly reduce the memory usage of a trained model by simultaneously improving the computational efficiency of the forward propagation process. While in the classical backpropagation framework commonly a $L^{1}$-norm penalty is considered to induce sparsity, we demonstrate that our optimal control-based technique applies also to $L^{0}$-’norm’ regularized machine learning problems. Thereby we are able to determine CNN models with an increased level of sparsity in the trained parameters compared to the $L^{1}$-based Elastic-Net approach [15], by simultaneously achieving a similar level of accuracy.
@@
In the next section, we introduce the interpretation of supervised learning tasks with CNN constraints as a discrete-time optimal control problem which allows us to formulate first-order optimality conditions in the form of a discrete Pontryagin maximum principle (PMP). The discrete PMP, as formulated by Halkin [16], is an extension of the work by Pontryagin et al. [17], who proved the PMP in the continuous-time optimal control framework; see also [18]. Opposed to the classical machine learning setting, where the first-order optimality conditions are based on the conditions of Karush-Kuhn-Tucker (KKT); see, for example, [19, 20], the PMP states that optimal trainable parameters are characterized by maximizing an auxiliary function at each layer, which is called the Hamilton-Pontryagin (HP) function. This maximization property allows the formulation of optimality conditions in the supervised learning framework even without differentiability assumptions on the loss function with respect to the trainable parameters. In Section 3, we formulate and discuss requirements for our PMP-based learning approach in the classical supervised learning framework, which is followed by a discussion of the discrete PMP for machine learning with CNNs in Section 4. In Section 5, we reason that the PMP motivates the formulation of a generalized backpropagation algorithm for CNNs based on a discrete version of the method of successive approximations (MSA). This early method was initially introduced by Krylov and Chernous’ko [21] for continuous-time optimal control problems, see also [22] and has since then undergone further improvements and extensions [12, 13, 14, 23, 24, 25]. In contrast to the original MSA method, these more recent modifications use an augmented HP function with a weighted penalization term to accomplish a stable optimization process. In particular, the choice of this augmentation weight has a decisive influence on the convergence of the algorithms, which makes methods such as the sequential quadratic Hamiltonian (SQH) scheme [13] particularly suitable for the definition of a stable training procedure, since it automatically determines an augmentation weight which guarantees a monotonic decrease in the loss functional value. We present a novel batch SQH method (bSQH), which enables the training of CNN architectures efficiently in a mini-batch fashion and improves upon the process of estimating the required augmentation weight incorporated into the augmented HP function. A decisive advantage of our methodology compared to the standard backpropagation approach lies in the formulation of the updating step which does not require the existence of gradients with respect to the trainable parameters. Consequently, our framework can accommodate supervised learning even with non-continuous regularizers and also enables the training of neural networks where the set of admissible parameters is discrete [14, 26, 27].
@@
We present in Section 6 a reformulation of non-residual CNN architectures such as LeNet-5 that enables the derivation of an analytic expression for the trainable parameters maximizing the augmented HP functions at each iteration. This procedure enables us to establish the bSQH scheme as a suitable algorithm for training sparse CNNs with a $L^{0}$-’norm’ based regularization technique. In Section 7, we prove the minimizing property of the bSQH algorithm by introducing a sufficient decrease condition for the mini-batch loss. Moreover, in addition to considerably extending our initial results from [13] to more general network structures and sparse learning techniques, we provide a novel convergence result. In particular, this result guarantees convergence of the bSQH scheme to trainable parameters satisfying the necessary optimality conditions of the discrete PMP. We validate the feasibility of the necessary conditions for convergence in Section 8. Further, in this section, we present numerical experiments of sparse network training in the framework of image classification with CNNs and an application to medical imaging. We numerically investigate a novel estimation strategy for the augmentation weights of the augmented HP function which allows a considerable reduction of the number of necessary inference steps involved in the training process. In this setting, we also compare the bSQH method in its full-batch setting with the mini-batch variant and discuss the computational advantages of the mini-batch approach. To conclude our numerical investigation, our PMP-based training procedure is successfully applied to solve a $L^{0}$-regularized supervised learning problem, where the involved CNN uses ReLU activations, batch normalization and max-pooling layers to classify abdominal CT images. A section of conclusion completes this work.
@@ 2 The supervised learning problem for CNNs
In the standard supervised learning setting, we have a data set composed of data-label pairs ${\cal D}:=\mathcal{X}\times\mathcal{Y},$ where $\mathcal{X}\subset{\mathbb{R}}^{n\times c}$ is a bounded set of input data and $\mathcal{Y}\subset{\mathbb{R}}^{m}$ is the corresponding set of labels. When considering the typical image classification task, each $x^{s}\in\mathcal{X}$ is interpreted as a suitably vectorized image, such that the dimension $n$ is the total number of pixels in each of the $c$ channels. A standard coloured image for example has three channels each dedicated to one of the primary colours, red, green and blue. While the number of channels is of importance in practice for our theoretical investigation, we can assume that $c=1$ holds without loss of generality using a suitable vectorization. Each $y^{s}\in\mathcal{Y}$ is a one-hot encoded class label representing the affiliation of the image $x^{s}$ to one of the $m$ different classes of images existing in $\mathcal{X}$. In supervised learning, one supposes that there exists an unknown, well-defined and continuous function $\mathcal{M}:\mathcal{X}\rightarrow\mathcal{Y}$ describing the data-label relation of elements in ${\cal D}$. Now, based on the information encoded in the data set, the supervised learning task is to find a suitable set of parameters ${\mathbf{u}}=(u_{l})_{l=1}^{L}$, where each $u_{l}$ is the element of an appropriate closed Euclidean subset $U_{l}$, such that the parametric, composite function
(1) ${\cal N}_{{\mathbf{u}}}:{\cal X}\to{\cal Y},\quad x\mapsto{\cal N}_{{\mathbf{u}}}(x)=\lambda_{L}(\;\cdot\;,u_{L})\circ\lambda_{L-1}(\;\cdot\;,u_{L-1})\circ\ldots\circ\lambda_{1}(x,u_{1}),$
approximates the mapping $\mathcal{M}$ up to a satisfying degree. The composite structure (1) itself is called the neural network and consists of $L\in{\mathbb{N}}$ layers $\lambda_{l}:{\mathbb{R}}^{n_{l}}\times U_{l}\rightarrow{\mathbb{R}}^{n_{l+1}}$ which, in the framework of CNNs, take the following form
@@
In (2), the transformation ${\cal S}_{l}:{\mathbb{R}}^{n_{l}}\rightarrow{\mathbb{R}}^{\tilde{n}_{l}}$ defines a projection or identity mapping for modelling skip connections and ${\cal P}_{l}:{\mathbb{R}}^{\tilde{n}_{l}}\rightarrow{\mathbb{R}}^{n_{l+1}}$ describes the process of a pooling operation or batch normalization [19]. These Pooling operations resemble the application of non-trainable filters like maximum or average filters and serve the purpose of reducing the spatial dimension of the processed data and at the same time enhance the network’s ability to extract important features. Based upon the universal approximation theorem of George Cybenko [28], the network’s core is represented by the non-linearity $\sigma_{l}({\cal W}_{l}(u)x+{\cal B}_{l}u)$, where for each layer $l=0,\ldots,L-1$, the transformation $\sigma_{l}:{\mathbb{R}}^{\tilde{n}_{l}}\rightarrow{\mathbb{R}}^{\tilde{n}_{l}}$, represents a non-linear component-wise defined activation function. To model the required linear affine transformation of the states, we introduce in the following the linear transformations ${\cal W}_{l}:U_{l}\rightarrow{\mathbb{R}}^{\tilde{n}_{l}\times n_{l}}$ and ${\cal B}_{l}:U_{l}\rightarrow{\mathbb{R}}^{\tilde{n}_{l}}$, which determine the layer’s type. A CNN is usually constructed using two types of layers; see also Figure 1: That is convolution layers for feature extraction and fully connected layers for classification of the extracted features. By taking into account that discrete convolutions can be reformulated as matrix multiplication [19], for each convolution layer, the transformation ${\cal W}_{l}$ resembles the form of a suitable sparse Toeplitz matrix. In contrast, this matrix is dense for a fully connected layer, where each element serves the purpose of an independently adjustable control.
@@
$\min_{{\mathbf{u}}\in U_{ad}}J({\mathbf{u}})$ $:=\mathbb{E}_{x\sim\mu}[\phi({\cal M}(x),{\cal N}_{{\mathbf{u}}}(x))]+\rho\;{\cal R}({\mathbf{u}}),$
@@
The function ${\cal R}:U_{ad}\rightarrow{\mathbb{R}}$ describes a regularizer which is weighted using the fixed scaling factor $\rho\geq 0$ and is applied in order to improve the stability of the model (1) or impose additional features such as sparsity of the parameters. Further, a suitable loss function $\phi:{\cal Y}\times{\cal Y}\rightarrow{\mathbb{R}}$ is defined which attains it’s minimum if all of its arguments coincide, i.e. ${\cal M}(x)={\cal N}_{{\mathbf{u}}}(x)$ for all $x\in{\cal X}$. However, in practice, only a finite number of observations is available in the form of a data-batch $B\subset{\cal D}$, so that the true probability distribution $\mu$ is unknown and only a finite set of evaluations of ${\cal M}$ exists. Under the assumption that every datapoint in $B$ is independent and identically distributed (i.i.d), the law of large numbers motivates to approximate $\mu$ using an empirical measure, which yields the so-called empirical risk minimization problem:
@@
where ${\cal L}_{B}({\mathbf{u}}):=\frac{1}{|B|}\sum_{s\in I_{B}}(\Phi_{s}\circ{\cal N}_{{\mathbf{u}}})(x^{s})$. For each $s\in I_{B}$, we define $\Phi_{s}:{\cal Y}\rightarrow{\mathbb{R}},\;y\mapsto\phi(y^{s},y),$ where $ys={\cal M}(x^{s})\in{\cal Y}$ denotes the label corresponding to a specific sample $x^{s}\in B$ with corresponding index $s\in I_{B}$.

### v2

@@ Abstract
A technique for determining this augmentation parameter is proposed, and the loss-reduction and convergence properties of the bSQH algorithm are analysed theoretically and validated numerically. Results of numerical experiments in the context of image classification with a sparsity-enforcing $L^{0}$-based regularizer demonstrate the effectiveness of the proposed method in full-batch and mini-batch modes.
@@ 1 Introduction
In the last decades, convolutional neural networks (CNNs) have become a standard choice for computer vision tasks and, in particular, medical imaging analysis, where they are applied to aid in detecting and diagnosing diseases and classifying tumour types based on MRI, CT or X-ray scans. Compared to its predecessors, like the multilayer perceptron (MLP), CNNs utilize convolutional layers with trainable filters in addition to fully connected layers, making them especially suited for processing data with a grid-like topology. This network architecture enables a more efficient extraction of local patterns and relevant features from an input image, by simultaneously requiring fewer parameters than fully connected architectures. The development of CNNs started in the 1980s when Kunihiko Fukushima [17] introduced the ’Neocognitron’ as the earliest precursor of convolutional networks. Almost one decade later, Yann LeCun [27] developed the LeNet-5 architecture for document recognition, representing a milestone in the design of modern CNNs. Over the last decades, a variety of architectures have contributed to the success of these network types, amongst which there is AlexNet [25], which popularized the use of ReLU activation functions and VGGNet [31] where an increased network depth was used, showing the benefits of using smaller filter sizes. However, it occurs that very deep architectures often face a vanishing gradient problem, preventing very deep networks from being trained efficiently using gradient-based training methods and backpropagation. To mitigate this vanishing gradient problem, Kaiming He et al. [21] developed the ResNet, a network architecture that uses skip connections and thereby enables the training of even deeper neural networks.
@@
Apart from this, the introduction of ResNet-like architectures gave birth to the dynamical system viewpoint of deep learning, where networks with a residual structure are related to numerical schemes for solving differential equations; see [15, 19, 8, 9, 32, 36]. The dynamical systems viewpoint offers important insights into the stability of a network’s forward propagation process by allowing the study of very deep networks from a continuous-time dynamical systems perspective. More importantly for our investigation, it motivates the development of novel learning techniques based on optimal control theory [28, 22].
@@
While serving as a motivation for developing new training algorithms, the continuous-time dynamical systems interpretation can be limiting, particularly when considering network architectures without skip connections, such as LeNet-5 and AlexNet. In those cases, the continuous-time interpretation becomes even more difficult when taking into account that CNN models use dimensionality reduction as a design approach for feature extraction. However, even if the relation to the continuous-time setting is lost for most modern network architectures, it has been shown that special optimal control-based techniques can still be applied to the machine learning framework, almost independently of the network’s structure [29]. Based thereon, in this work, we considerably extend our optimal control-based learning technique presented in [22] to the framework of CNNs. For this purpose, we construct a CNN as a discrete-time controlled system including trainable filters, dimensionality reduction in the form of pooling operations, and a sequence of fully-connected layers for classification. Further, for this extension and within this special structure, we also take into account the computational complexity arising from the increased number of parameters and dimensionality of data, which can impose memory constraints. To address these challenges, we derive a novel mini-batch variant of our optimal control-based learning algorithm. Furthermore, we include sparse network training in the discussion of our algorithm, where the aim is to determine the essential parameters for a given model by forcing irrelevant or redundant ones to be zero.
@@
In the context of CNNs, where models are often over-parametrized, deriving sparse solutions can significantly reduce the memory usage of a trained model while improving the computational efficiency of the forward propagation process. In the classical backpropagation framework, commonly a $L^{1}$-norm penalty is considered to induce sparsity. However, we demonstrate that with our optimal control-based technique, also $L^{0}$-’norm’ regularized machine learning problems can be solved. In this way, we can determine CNN models with a higher level of sparsity in the trained parameters compared to the $L^{1}$-based Elastic-Net approach [45], while simultaneously achieving a similar level of accuracy.
@@
In the next section, we introduce the interpretation of supervised learning tasks with CNN constraints as a discrete-time optimal control problem, which enables us to formulate first-order optimality conditions in the form of a discrete Pontryagin maximum principle (PMP). The discrete PMP, as formulated by Halkin [20], is an extension of the work by Pontryagin et al. [34], who proved the PMP in the continuous-time optimal control framework; see also [4]. Opposed to the classical machine learning setting, where the first-order optimality conditions are based on the conditions of Karush-Kuhn-Tucker (KKT); see, for example, [1, 33], the PMP states that optimal trainable parameters are characterized by maximizing an auxiliary function at each layer, which is called the Hamilton-Pontryagin (HP) function. This maximization property allows the formulation of optimality conditions in the supervised learning framework even without differentiability assumptions on the loss function with respect to the trainable parameters. In Section 3, we formulate and discuss requirements for our PMP-based learning approach in the classical supervised learning framework, which is followed by a discussion of the discrete PMP for machine learning with CNNs in Section 4. In Section 5, we reason that the PMP motivates the formulation of a generalized backpropagation algorithm for CNNs based on a discrete version of the method of successive approximations (MSA). This early method was initially introduced by Krylov and Chernous’ko [26] for continuous-time optimal control problems, see also [24], and has since then undergone further improvements and extensions [28, 22, 29, 37, 10, 6]. In contrast to the original MSA method, these more recent modifications use an augmented HP function with a weighted penalization term to accomplish a stable optimization process. In particular, the choice of this augmentation weight has a decisive influence on the convergence of the algorithms, which makes methods such as the sequential quadratic Hamiltonian (SQH) scheme [22] particularly suitable for the definition of a stable training procedure, since it automatically determines an augmentation weight which guarantees a monotonic decrease in the loss functional value. We present a novel batch SQH method (bSQH), which enables efficient training of CNN architectures in a mini-batch fashion and improves upon the process of estimating the required augmentation weight incorporated into the augmented HP function. A decisive advantage of our methodology compared to the standard backpropagation approach lies in the formulation of the updating step, which does not require the existence of gradients with respect to the trainable parameters. Consequently, our framework can accommodate supervised learning even with non-continuous regularizers and also enables the training of neural networks where the set of admissible parameters is discrete [29, 2, 3].
@@
We present in Section 6 a reformulation of non-residual CNN architectures such as LeNet-5 that enables the derivation of an analytic expression for the trainable parameters maximizing the augmented HP functions at each iteration. This allows us to establish the bSQH scheme as a suitable algorithm for training sparse CNNs with a $L^{0}$-’norm’ based regularization technique. In Section 7, we prove the minimizing property of the bSQH algorithm by introducing a sufficient decrease condition for the mini-batch loss. Moreover, in addition to considerably extending our initial results from [22] to more general network structures and sparse learning techniques, we provide a novel convergence result. In particular, this result guarantees convergence of the bSQH scheme to trainable parameters satisfying the necessary optimality conditions of the discrete PMP. We validate the feasibility of the necessary conditions for convergence in Section 8. Further, in this section, we present numerical experiments of sparse network training in the framework of image classification with CNNs and an application to medical imaging. We numerically investigate a novel estimation strategy for the augmentation weights of the augmented HP function, which leads to a considerable reduction in the number of necessary inference steps required for the training process. In this setting, we also compare the bSQH method in its full-batch setting with the mini-batch variant and discuss the computational advantages of the mini-batch approach. To conclude our numerical investigation, our PMP-based training procedure is successfully applied to solve a $L^{0}$-regularized supervised learning problem, where the involved CNN uses ReLU activations, batch normalization and max-pooling layers to classify abdominal CT images. A section of conclusion completes this work.
@@ 2 The supervised learning problem for CNNs
In the standard supervised learning setting, we have a dataset composed of data-label pairs ${\cal D}:=\mathcal{X}\times\mathcal{Y},$ where $\mathcal{X}\subset{\mathbb{R}}^{n\times c}$ is a bounded set of input data and $\mathcal{Y}\subset{\mathbb{R}}^{m}$ is the corresponding set of labels. When considering the typical image classification task, each $x^{s}\in\mathcal{X}$ is interpreted as a suitably preprocessed and vectorized image, where the dimension $n$ is the total number of pixels in each of the $c$ channels. A standard coloured image, for example, has three channels, each dedicated to one of the primary colours, red, green and blue. For our theoretical investigation, we can assume without loss of generality that $c=1$ holds by supposing a suitable vectorization. Each $y^{s}\in\mathcal{Y}$ is a one-hot encoded label representing the affiliation of the image $x^{s}$ to one of the $m$ different classes of images contained in $\mathcal{X}$. We suppose that there exists an unknown, well-defined and continuous mapping $\mathcal{M}:\mathcal{X}\rightarrow\mathcal{Y}$ describing the data-label relation of elements in $\cal D$}. Then, based on the information encoded in the dataset, the supervised learning task is to find a suitable set of parameters ${\mathbf{u}}=(u_{l})_{l=0}^{L-1}$, where each $u_{l}$ is the element of an appropriate closed Euclidean subset $U_{l}$, such that the following parametric, composite function approximates the mapping $\mathcal{M}$ up to a satisfying degree
(1) ${\cal N}_{{\mathbf{u}}}:{\cal X}\to{\cal Y},\quad x\mapsto{\cal N}_{{\mathbf{u}}}(x)=\lambda_{L-1}(\;\cdot\;,u_{L-1})\circ\lambda_{L-2}(\;\cdot\;,u_{L-2})\circ\ldots\circ\lambda_{0}(x,u_{0}).$
The composite structure (1) itself is called the neural network and consists of $L\in{\mathbb{N}}$ layers $\lambda_{l}:{\mathbb{R}}^{n_{l}}\times U_{l}\rightarrow{\mathbb{R}}^{n_{l+1}}$ which, in the framework of CNNs, take the following form
@@
In (2), the transformation ${\cal S}_{l}:{\mathbb{R}}^{n_{l}}\rightarrow{\mathbb{R}}^{\tilde{n}_{l}}$ defines a projection or identity mapping for modeling skip connections and ${\cal P}_{l}:{\mathbb{R}}^{\tilde{n}_{l}}\rightarrow{\mathbb{R}}^{n_{l+1}}$ describes the process of a pooling operation or batch normalization [1]. Pooling operations resemble the application of non-trainable filters, such as maximum or average filters and serve the purpose of reducing the spatial dimension of the processed data while enhancing the network’s ability to extract important features. Based upon the universal approximation theorem of George Cybenko [12], the network’s core is represented by the non-linearity $\sigma_{l}({\cal W}_{l}(u)x+{\cal B}_{l}u)$, where for each layer $l=0,\ldots,L-1$, the transformation $\sigma_{l}:{\mathbb{R}}^{\tilde{n}_{l}}\rightarrow{\mathbb{R}}^{\tilde{n}_{l}}$, is a non-linear component-wise defined activation function. To model the linear affine transformation of the states in (2), we introduce the linear transformations ${\cal W}_{l}:U_{l}\rightarrow{\mathbb{R}}^{\tilde{n}_{l}\times n_{l}}$ and ${\cal B}_{l}:U_{l}\rightarrow{\mathbb{R}}^{\tilde{n}_{l}}$ which determine the type of a layer. A CNN is commonly constructed with two types of layers: Convolutional layers for feature extraction and fully connected layers for classification of the extracted features; see Figure 1. By taking into account that discrete convolutions can be reformulated as matrix multiplication [1], in each convolutional layer, the transformation ${\cal W}_{l}(u)$ takes the form of a sparse Toeplitz matrix. In contrast, for fully connected layers, this transformation is represented by a dense connectivity matrix, where each element acts as an independently adjustable weight.
@@
$\min_{{\mathbf{u}}\in U_{ad}}J({\mathbf{u}})$ $=\mathbb{E}_{x\sim\mu}[\phi({\cal M}(x),{\cal N}_{{\mathbf{u}}}(x))]+\rho\;{\cal R}({\mathbf{u}}),$
@@
[…]
@@
where ${\cal L}_{B}({\mathbf{u}}):=\frac{1}{|B|}\sum_{s\in I_{B}}(\Phi_{s}\circ{\cal N}_{{\mathbf{u}}})(x^{s})$. For each $s\in I_{B}$, we define $\Phi_{s}:{\cal Y}\rightarrow{\mathbb{R}},\;y\mapsto\phi(y^{s},y),$ where $y^{s}={\cal M}(x^{s})\in{\cal Y}$ denotes the label corresponding to a specific sample $x^{s}\in B$ with corresponding index $s\in I_{B}$.